# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ Abstract
There has been an increasing interest on summary–free versions of approximate Bayesian computation (abc), which replace distances among summaries with discrepancies between the empirical distributions of the observed data and the synthetic samples generated under the proposed parameter values. The success of these solutions has motivated theoretical studies on the limiting properties of the induced posteriors. However, current results (i) are often tailored to a specific discrepancy, (ii) require, either explicitly or implicitly, regularity conditions on both the data generating process and the assumed statistical model, and (iii) yield concentration bounds depending on sequences of control functions that are not made explicit. As such, there is still the lack of a theoretical framework for summary–free abc that (i) is unified, instead of discrepancy–specific, (ii) does not necessarily require to constrain the analysis to data generating processes and statistical models meeting specific regularity conditions, but rather facilitates the derivation of limiting properties that hold uniformly, and (iii) relies on verifiable assumptions that provide more explicit concentration bounds clarifying which factors govern the limiting behavior of the abc posterior. We address this gap via a novel theoretical framework that introduces the concept of Rademacher complexity in the analysis of the limiting properties for discrepancy–based abc posteriors, including in non–i.i.d. and misspecified settings. This yields a unified theory that relies on constructive arguments and provides more informative asymptotic results and uniform concentration bounds, even in settings not covered by current studies. These advancements are obtained by relating the limiting properties of summary–free abc posteriors to the behavior of the Rademacher complexity associated with the chosen discrepancy within the family of integral probability semimetrics (ips). The ips class extends summary–based abc, and also includes the widely–implemented Wasserstein distance and maximum mean discrepancy (mmd), among others. As clarified through a focus of these results on the mmd case and via illustrative simulations, this novel theoretical perspective yields an improved understanding of summary–free abc.
@@ 1 Introduction
The increasing complexity of statistical models in modern applications not only yields intractable likelihoods, but also raises substantial challenges in the identification of effective summary statistics (e.g., Fearnhead & Prangle, 2012; Marin et al., 2014; Frazier et al., 2018). Such drawbacks have motivated an increasing adoption of abc solutions, along with a shift away from summary–based implementations (e.g., Marin et al., 2012) and towards summary–free strategies relying on discrepancies among the empirical distributions of the observed and synthetic data (e.g., Drovandi & Frazier, 2022). These solutions only require that simulating from the model is feasible, and provide samples from an approximate posterior distribution for the parameter of interest ${\theta\in\Theta\subseteq\mathbb{R}^{p}}$. This is achieved by retaining all those values of $\theta$, drawn from the prior, which produced synthetic samples $z_{1:m}=(z_{1},\ldots,z_{m})$ from the model $\mu_{\theta}$ whose empirical distribution is sufficiently close to the one of the observed data $y_{1:n}=(y_{1},\ldots,y_{n})$, under the chosen discrepancy.
@@
Notable examples of the above implementations are abc versions that employ maximum mean discrepancy (mmd) (Park et al., 2016), Kullback–Leibler (kl) divergence (Jiang et al., 2018), Wasserstein distance (Bernton et al., 2019), energy statistic (Nguyen et al., 2020), the Hellinger and Cramer–von Mises distances (Frazier, 2020), and the $\gamma$–divergence (Fujisawa et al., 2021); see also Gutmann et al. (2018), Forbes et al. (2021) and Wang et al. (2022) for additional examples of summary–free abc strategies. By overcoming the need to preselect summaries, all these solutions reduce the potential information loss of summary–based abc, thus yielding improved performance in simulation studies and illustrative applications. These promising empirical results have motivated active research on the theoretical properties of the induced abc posterior, with a main focus on the limiting behavior under different asymptotic regimes for the tolerance threshold and the sample size (Jiang et al., 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Fujisawa et al., 2021). Among these regimes, of particular interest are the two situations in which the abc threshold is either fixed or progressively shrinks as both $n$ and $m$ diverge. In the former case, the theory focus is on assessing whether the control on the discrepancy among the empirical distributions established by the selected abc threshold yields a pseudo–posterior whose asymptotic form guarantees the same threshold–control on the discrepancy among the underlying truths (e.g., Jiang et al., 2018). The latter addresses instead the more challenging theoretical question on whether a suitably–decaying abc threshold can provide meaningful rates of concentration for the sequence of abc posteriors around those $\theta$ values yielding a $\mu_{\theta}$ close to the data generating process $\mu^{*}$, as $n$ and $m$ diverge (e.g., Bernton et al., 2019).
@@
Available contributions along these lines of research have the merit of providing theoretical support to several important versions of summary–free abc. However, current theory is often tailored to the specific discrepancy analyzed, and generally relies on difficult–to–verify existence assumptions and concentration inequalities that constrain the analysis, either implicitly or explicitly, to data generating processes and statistical models which meet suitable regularity conditions, thereby lacking results that hold uniformly. Recalling Bernton et al. (2019) and Nguyen et al. (2020), this yields also concentration bounds involving sequences of control functions which are not made explicit. As such, although convergence and concentration can still be proved, the core factors that govern these limiting properties remain yet unexplored, thus limiting the methodological impact of current theory, while hindering the derivation of novel informative results in more challenging settings. For example, the available theoretical studies pre–assume that the discrepancy among the empirical distributions $\hat{\mu}_{z_{1:m}}$ and $\hat{\mu}_{y_{1:n}}$ suitably converges to the one among the corresponding truths $\mu_{\theta}$ and $\mu^{*}$, or, alternatively, that both $\hat{\mu}_{z_{1:m}}$ and $\hat{\mu}_{y_{1:n}}$ converge, under the selected discrepancy, to $\mu_{\theta}$ and $\mu^{*}$, respectively. While these assumptions can be verified under suitable conditions and for specific discrepancies, as clarified within Sections 2 and 3, an in–depth theoretical understanding of summary–free abc necessarily requires relating convergence to the learning properties of the selected discrepancy, rather than pre–assuming it. Such a more precise but yet–unexplored theoretical treatment has the potentials to shed light on the factors that govern the limiting behavior of discrepancy–based abc posteriors. In addition, it could possibly provide more general, verifiable and explicit sufficient conditions under which popular discrepancies are guaranteed to ensure that the previously pre–assumed convergence and concentration hold uniformly over models and data generating processes, while facilitating the study of the limiting behavior of discrepancy–based abc posteriors in more general situations where these convergence guarantees may lack.
@@
In this article, we address the aforementioned endeavor by introducing an innovative theoretical framework which analyses the limiting properties of discrepancy–based abc posteriors, under a unified perspective and for different asymptotic regimes, through the concept of Rademacher complexity (e.g., Wainwright, 2019, Chapter 4), within the general class of integral probability semimetrics (ips) (see e.g., Müller, 1997; Sriperumbudur et al., 2012). This class naturally generalizes summary–based abc and includes the widely–implemented mmd and Wasserstein distance, among others. As clarified in Sections 2, 3 and 6, such a perspective, which to the best of our knowledge is novel in abc, allows to derive unified, informative and uniform concentration bounds for abc posteriors under several discrepancies, in possibly misspecified and non–i.i.d. contexts. Moreover, it relies on more constructive arguments that clarify under which sufficient conditions a discrepancy within the ips class guarantees uniform convergence and concentration of the induced abc posterior; i.e., without necessarily requiring suitable regularity conditions for the underlying data generating process $\mu^{*}$ and the assumed statistical model. This yields an important theoretical and methodological advancement, provided that $\mu^{*}$ is often unknown in practice and, hence, verifying regularity conditions on the data generating process is generally unfeasible.
@@
Crucially, the theoretical framework we introduce allows to prove informative theoretical results even in yet–unexplored settings that possibly lack those convergence guarantees assumed in available literature. More specifically, in these settings we derive novel upper and lower bounds for the limiting acceptance probabilities that clarify in which contexts the abc posterior is still well–defined for $n$ large enough. When this is the case, it is further possible to obtain informative supersets for the support of such a posterior. These show that, when relaxing standard convergence assumptions employed in state–of–the–art theoretical studies, the control established by a fixed abc threshold on the discrepancy among the empirical distributions does not necessarily translate, asymptotically, into the same control on the discrepancy among the corresponding truths, but rather yields an upper bound equal to the sum between the abc threshold and a multiple of the Rademacher complexity, namely a measure of richness of the class of functions that uniquely identify the chosen ips; see Section 3.1.
@@
The above results clarify the fundamental relation among the limiting behavior of abc posteriors and the learning properties of the chosen discrepancy, when measured via Rademacher complexity. Moreover, the bounds derived clarify that a sufficient condition to recover a limiting pseudo–posterior with the same threshold–control on the discrepancy among the truths as the one enforced on the corresponding empirical distributions, is that the selected discrepancy has a Rademacher complexity vanishing to zero in the large–data limit. As proved in Section 3.2, this setting also allows constructive derivations of novel, informative and uniform concentration bounds for discrepancy–based abc posteriors in the challenging regime where the threshold shrinks towards zero as both $m$ and $n$ diverge. This is facilitated by the existence of meaningful upper bounds for the Rademacher complexity of popular abc discrepancies, along with the availability of constructive conditions for the derivation of these bounds (e.g., Sriperumbudur et al., 2012) which leverage fundamental connections among such a complexity measure and other key quantities in statistical learning theory, such as the Vapnik–Chervonenkis (vc) dimension and the notion of uniform Glivenko–Cantelli classes (see e.g., Wainwright, 2019, Chapter 4). This yields an improved understanding of the factors that govern the concentration of discrepancy–based abc posteriors under a unified perspective that further allows to (i) quantify rates of concentration and (ii) directly translate any advancement on Rademacher complexity into novel abc theory. The former advantage is illustrated within Section 4 through a specific focus on mmd with routinely–implemented bounded and unbounded kernels, whereas the latter is clarified in Section 6, where we extend the theory from Section 3 to non–i.i.d. settings, leveraging results in Mohri & Rostamizadeh (2008) on Rademacher complexity under $\beta$–mixing processes (e.g., Doukhan, 1994).
@@
The illustrative simulation studies within Sections 5 and 7 show that the theoretical results derived in Sections 3, 4 and 6 find empirical evidence in practice, including also in scenarios based on model misspecification, data contamination and non–i.i.d. data generating processes. These findings suggest that discrepancies with guarantees of uniformly–vanishing Rademacher complexity provide a robust and sensible choice when the assumed statistical model and/or the underlying data generating process do not necessarily meet specific regularity conditions, or it is not possible to verify these conditions. This is a common situation in applications, provided that the data generating process is often unknown in practice.
@@
As discussed in Section 8, the unexplored bridge between discrepancy–based abc and the Rademacher complexity introduced in this article can also be leveraged to derive even more general theory by exploiting the active literature on the Rademacher complexity. For example, combining our novel perspective with the recently–derived unified treatment of ips and $f$–divergences (Agrawal and Horel, 2021; Birrell et al., 2022) might set the premises to derive similarly–tractable, interpretable and general results for other discrepancies employed within abc, such as the Kullback–Leibler divergence (Jiang et al., 2018) and Hellinger distance (Frazier, 2020). More generally, our contribution can have direct implications even beyond abc, in particular on generalized Bayesian inference via pseudo–posteriors based on discrepancies (e.g., Bissiri et al., 2016; Chérief-Abdellatif & Alquier, 2020; Matsubara et al., 2022). The proofs of the theoretical results can be found in the Supplementary Material.
2 Integral probability semimetrics (ips) and Rademacher complexity
@@
Let $y_{1:n}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$ denote an i.i.d. sample from $\mu^{*}\in\mathcal{P}(\mathcal{Y})$, where $\mathcal{P}(\mathcal{Y})$ is the space of probability measures on $\mathcal{Y}$, and assume that $\mathcal{Y}$ is a metric space endowed with the distance $\rho$; see Section 6 for extensions to non–i.i.d. settings. Given a statistical model $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ within $\mathcal{P}(\mathcal{Y})$ and a prior distribution $\pi$ on $\theta$, rejection abc iteratively samples $\theta$ from $\pi$, draws a synthetic i.i.d. sample $z_{1:m}=(z_{1},\ldots,z_{m})$ from $\mu_{\theta}$, and retains $\theta$ as a sample from the abc posterior $\pi_{n}^{(\varepsilon_{n})}$ if the discrepancy $\Delta(z_{1:m},y_{1:n})$ between the synthetic and the observed data is below a preselected threshold $\varepsilon_{n}\geq 0$. Although $m$ may differ from $n$, here we follow common practice in abc theory (Bernton et al., 2019; Frazier, 2020) and set ${m=n}$. Rejection abc does not sample from the exact posterior distribution $\pi_{n}(\theta)\propto\pi(\theta)\mu_{\theta}^{n}(y_{1:n}),$ but rather from the abc posterior
@@
whose properties clearly depend on the chosen discrepancy $\Delta(\cdot,\cdot)$. Within summary–based abc, $\Delta(z_{1:n},y_{1:n})$ is a suitable distance, typically Euclidean, among summaries computed from the synthetic and observed data. However, recalling, e.g., Marin et al. (2014) and Frazier et al. (2018), the identification of summaries that do not lead to information loss is challenging for those complex models requiring abc implementations.
@@
where $\delta_{x}$ denotes the Dirac delta at a generic $x\in\mathcal{Y}$. Popular examples are abc versions based on mmd, kl divergence, Wasserstein distance, energy statistic, Hellinger and Cramer–von Mises distances, and $\gamma$–divergence, whose limiting properties have been studied in Park et al. (2016), Jiang et al. (2018), Bernton et al. (2019), Nguyen et al. (2020), Frazier (2020) and Fujisawa et al. (2021) under different asymptotic regimes and with a common reliance on existence assumptions to ease the proofs. As a first step toward our unified and constructive theoretical framework, we shall emphasize that, although most of the above contributions treat discrepancies separately, some of these choices share, in fact, a common origin. For example, mmd, Wasserstein distance and the energy statistic all belong to the ips class (Müller, 1997) in Definition 2.1. Such a class, also includes summary–based distances.
Definition 2.1 (Integral probability semimetric – ips).
Let $(\mathcal{Y},\mathscr{A})$ denote a measure space, $g:{\mathcal{Y}\to[1,\infty)}$ a measurable function, and $\mathfrak{B}_{g}$ the set of all measurable functions $f:\mathcal{Y}\to\mathds{R}$ such that $||f||_{g}:=\sup_{y\in\mathcal{Y}}|f(y)|/g(y)<\infty.$ Then, for any $\mathfrak{F}\subseteq\mathfrak{B}_{g}$, a generic integral probability semimetric $\mathcal{D}_{\mathfrak{F}}$ between $\mu_{1},\mu_{2}\in\mathcal{P}(\mathcal{Y})$ is defined as
@@
where $\sup\nolimits_{f\in\mathfrak{F}}$ denotes the $\sup$ over the prespecified class of functions $\mathfrak{F}$.
Examples 2.12.3 show that routinely–employed discrepancies both in summary–free and summary–based abc (Park et al., 2016; Bernton et al., 2019; Nguyen et al., 2020; Drovandi & Frazier, 2022) are, in fact, ips with a known characterizing family $\mathfrak{F}$ that uniquely identifies each discrepancy.
@@
(Wasserstein distance). Let us consider the Lipschitz seminorm defined as $||f||_{L}{:=}\sup\{|f(x)-f(x^{\prime})|/\rho(x,x^{\prime}):x\neq x^{\prime}\mbox{ in }\mathcal{Y}\}$. If $\mathfrak{F}=\{f:||f||_{L}\leq 1\}$, then $\mathcal{D}_{\mathfrak{F}}$ coincides with the Kantorovich metric. Recalling the Kantorovich–Rubinstein theorem (e.g., Dudley, 2018), when $\mathcal{Y}$ is separable this metric coincides with the dual representation of the Wasserstein distance, which is therefore recovered as an example of ips. Recall also that such a distance is defined on the space of probability measures having finite first–order moments (e.g., Sriperumbudur et al., 2012; Bernton et al., 2019).
@@
(mmd and energy distance). Given a positive–definite kernel $k(\cdot,\cdot)$ on $\mathcal{Y}\times\mathcal{Y}$, let $\mathfrak{F}=\{f:||f||_{\mathcal{H}}\leq 1\}$, where $\mathcal{H}$ is the reproducing kernel Hilbert space corresponding to $k(\cdot,\cdot)$. Then $\mathcal{D}_{\mathfrak{F}}$ is the mmd (e.g., Muandet et al., 2017). When $k(\cdot,\cdot)$ is characteristic, then $\mathcal{D}_{\mathfrak{F}}$ is not only a semimetric, but also a metric — i.e., $\mathcal{D}_{\mathfrak{F}}(\mu_{1},\mu_{2})=0$ implies $\mu_{1}=\mu_{2}$. Relevant examples of routinely–implemented characteristic kernels in $\mathcal{Y}=\mathbb{R}^{d}$ are the Gaussian $\exp(-\|x-x^{\prime}\|^{2}/\sigma^{2})$ and Laplace $\exp(-\|x-x^{\prime}\|/\sigma)$ ones (Muandet et al., 2017). Note that mmd is also directly related to energy distance, due to the correspondence among positive–definite kernels and negative–definite functions (Sejdinovic et al., 2013).

### v5

# Concentration of discrepancy-based approximate Bayesian computation via Rademacher complexity

@@ Abstract
There has been an increasing interest on summary-free solutions for approximate Bayesian computation (abc) which replace distances among summaries with discrepancies between the empirical distributions of the observed data and the synthetic samples generated under the proposed parameter values. The success of these strategies has motivated theoretical studies on the limiting properties of the induced posteriors. However, there is still the lack of a theoretical framework for summary-free abc that (i) is unified, instead of discrepancy-specific, (ii) does not necessarily require to constrain the analysis to data generating processes and statistical models meeting specific regularity conditions, but rather facilitates the derivation of limiting properties that hold uniformly, and (iii) relies on verifiable assumptions that provide more explicit concentration bounds clarifying which factors govern the limiting behavior of the abc posterior. We address this gap via a novel theoretical framework that introduces the concept of Rademacher complexity in the analysis of the limiting properties for discrepancy-based abc posteriors, including in non-i.i.d. and misspecified settings. This yields a unified theory that relies on constructive arguments and provides more informative asymptotic results and uniform concentration bounds, even in settings not covered by current studies. These advancements are obtained by relating the asymptotic properties of summary-free abc posteriors to the behavior of the Rademacher complexity associated with the chosen discrepancy within the family of integral probability semimetrics (ips). The ips class extends summary-based distances, and also includes the widely-implemented Wasserstein distance and maximum mean discrepancy (mmd), among others. As clarified in specialized theoretical analyses of popular ips discrepancies and via illustrative simulations, this new perspective improves the understanding of summary-free abc.
62F15,
@@ 1 Introduction
The growing complexity of statistical models in modern applications not only yields intractable likelihoods, but also raises substantial challenges in the identification of effective summary statistics (see e.g., Fearnhead and Prangle, 2012; Marin et al., 2014; Frazier et al., 2018). Such drawbacks have motivated an increasing adoption of abc solutions, along with a shift away from summary-based implementations (see e.g., Marin et al., 2012) and towards summary-free strategies relying on discrepancies among the empirical distributions of the observed and synthetic data (see e.g., Drovandi and Frazier, 2022). These solutions provide samples from an approximate posterior distribution for the parameter of interest ${\theta\in\Theta\subseteq\mathbb{R}^{p}}$ under the only requirement that simulating from the assumed model $\mu_{\theta}$ is feasible. This is achieved by retaining all those values of $\theta$, drawn from the prior, that produced synthetic samples $z_{1:m}=(z_{1},\ldots,z_{m})$ from $\mu_{\theta}$ whose empirical distribution is close enough to the one of the observed data $y_{1:n}=(y_{1},\ldots,y_{n})$, under the chosen discrepancy.
@@
Remarkable examples of the above implementations are abc versions that employ maximum mean discrepancy (mmd) (Park, Jitkrittum and Sejdinovic, 2016), Kullback–Leibler (kl) divergence (Jiang, Wu and Wong, 2018), Wasserstein distance (Bernton et al., 2019), energy statistic (Nguyen et al., 2020), Hellinger and Cramer–von Mises distances (Frazier, 2020), and $\gamma$-divergence (Fujisawa et al., 2021); see also Gutmann et al. (2018), Forbes et al. (2021) and Wang, Kaji and Rockova (2022) for additional examples of summary-free abc strategies. By overcoming the need to pre-select summaries, all these solutions reduce the potential information loss of summary-based abc, thereby yielding improved performance in simulation studies and illustrative applications. These promising empirical results have motivated active research on the theoretical properties of the induced abc posterior, with a main focus on the limiting behavior under different asymptotic regimes for the tolerance threshold and the sample size (Jiang, Wu and Wong, 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Fujisawa et al., 2021). Among these regimes, of particular interest are the two situations in which the abc threshold is either fixed or progressively shrinks as both $n$ and $m$ diverge. In the former case, the theory focus is on assessing whether the control on the discrepancy among the empirical distributions established by the selected abc threshold yields a pseudo-posterior whose asymptotic form guarantees the same threshold-control on the discrepancy among the underlying truths (e.g., Jiang, Wu and Wong, 2018). The latter addresses instead the more challenging theoretical question on whether a suitably-decaying abc threshold can provide meaningful rates of concentration for the sequence of abc posteriors around those $\theta$ values yielding a $\mu_{\theta}$ close enough to the data generating process $\mu^{*}$, as $n$ and $m$ diverge (e.g., Bernton et al., 2019).
@@
Available contributions along these lines of research have the merit of providing theoretical support to several versions of summary-free abc. However, current theory is often tailored to the specific discrepancy analyzed, and generally relies on difficult-to-verify existence assumptions and concentration inequalities that constrain the analysis, either implicitly or explicitly, to data generating processes and statistical models which meet suitable regularity conditions, thereby lacking results that hold uniformly. Recalling Bernton et al. (2019) and Nguyen et al. (2020), this also yields concentration bounds involving sequences of control functions which are not made explicit. As such, although convergence and concentration can still be proved, the core factors that govern these asymptotic properties remain yet unexplored, thus limiting the methodological impact of current theory, while hindering the derivation of novel informative results in more challenging settings. For example, the available theoretical studies pre-assume that the discrepancy among the empirical distributions $\hat{\mu}_{z_{1:m}}$ and $\hat{\mu}_{y_{1:n}}$ suitably converges to the one among the corresponding truths $\mu_{\theta}$ and $\mu^{*}$, or, alternatively, that both $\hat{\mu}_{z_{1:m}}$ and $\hat{\mu}_{y_{1:n}}$ converge, under the selected discrepancy, to $\mu_{\theta}$ and $\mu^{*}$, respectively. While these assumptions can be verified under suitable conditions and for specific discrepancies, as clarified within Sections 2 and 3, an in-depth theoretical understanding of summary-free abc necessarily requires relating convergence to the learning properties of the selected discrepancy, rather than pre-assuming it. Such a more precise but yet-unexplored theoretical treatment has the potentials to shed light on the factors that govern the limiting behavior of discrepancy-based abc posteriors. In addition, it could possibly provide more general, verifiable and explicit sufficient conditions under which popular discrepancies are guaranteed to ensure that the previously pre-assumed convergence and concentration hold uniformly over models and data generating processes, while facilitating the study of the limiting behavior of discrepancy-based abc posteriors in more general situations where these convergence guarantees may lack.
@@
In this article, we address the above gap by introducing an innovative theoretical framework which analyses the limiting properties of discrepancy-based abc posteriors, under a unified perspective and for different asymptotic regimes, through the concept of Rademacher complexity (see e.g., Wainwright, 2019, Chapter 4), within the general class of integral probability semimetrics (ips) (e.g., Müller, 1997; Sriperumbudur et al., 2012). Such a class naturally generalizes distances among summaries and includes the widely-implemented mmd and Wasserstein distance, among others. As clarified in Sections 23 and in Appendix C of the Supplementary Material, this perspective, which to the best of our knowledge is novel within abc, allows to derive unified, informative and uniform concentration bounds for abc posteriors under several discrepancies, in possibly misspecified and non-i.i.d. contexts. Moreover, it relies on more constructive arguments that clarify under which sufficient conditions a discrepancy within the ips class guarantees uniform convergence and concentration of the induced abc posterior; i.e., without necessarily requiring suitable regularity conditions for the underlying data generating process $\mu^{*}$ and the assumed statistical model. This yields an important theoretical and methodological advancement, since $\mu^{*}$ is often unknown in practice and, hence, verifying regularity conditions on the data generating process is generally unfeasible.
@@
Crucially, the theoretical framework we introduce allows to prove informative theoretical results even in yet-unexplored settings that possibly lack those convergence guarantees assumed in the literature. More specifically, in these settings we derive novel upper and lower bounds for the limiting acceptance probabilities that clarify in which contexts the abc posterior is still well-defined for $n$ large enough. When this is the case, it is further possible to obtain informative supersets for the support of such a posterior. These show that, when relaxing standard convergence assumptions employed in state-of-the-art theoretical studies, the control established by a fixed abc threshold on the discrepancy among the empirical distributions does not necessarily translate, asymptotically, into the same control on the discrepancy among the corresponding truths, but rather yields an upper bound equal to the sum between the abc threshold and a multiple of the Rademacher complexity, namely a measure of richness of the class of functions that uniquely identify the chosen ips; see Section 3.1.
@@
The above results clarify the fundamental relation among the limiting behavior of abc posteriors and the learning properties of the chosen discrepancy, when measured via Rademacher complexity. In addition, the bounds derived clarify that a sufficient condition to recover a limiting pseudo-posterior with the same threshold-control on the discrepancy among the truths as the one enforced on the corresponding empirical distributions, is that the selected discrepancy has a Rademacher complexity vanishing to zero in the large-data limit. As proved within Section 3.2, this setting also allows constructive derivations of novel, informative and uniform concentration bounds for discrepancy-based abc posteriors in the challenging regime where the threshold shrinks towards zero as both $m$ and $n$ diverge. This is facilitated by the existence of meaningful upper bounds for the Rademacher complexity associated to popular abc discrepancies, along with the availability of constructive conditions for the derivation of these bounds (Sriperumbudur et al., 2012). Such results leverage fundamental connections among the Rademacher complexity and other key quantities in statistical learning theory, such as the Vapnik–Chervonenkis (vc) dimension and the notion of uniform Glivenko–Cantelli classes (e.g., Wainwright, 2019, Chapter 4). This yields an improved understanding of the factors that govern the concentration of discrepancy-based abc posteriors under a unified perspective that further allows to (i) quantify rates of concentration and (ii) directly translate any advancement on Rademacher complexity into novel abc theory. Section 4 illustrates point (i) through a specific focus on mmd with routinely-implemented bounded kernels, and also clarifies that in the absence of guarantees on uniformly-vanishing Rademacher complexity (e.g., under Wasserstein distance in unbounded data spaces) concentration results can be still derived, but at the expense of regularity conditions for the data generating process $\mu^{*}$ and the assumed model. Point (ii) is clarified in Appendix C of the Supplementary Material, where we extend the theory from Section 3 to non-i.i.d. settings, leveraging results in Mohri and Rostamizadeh (2008) on the Rademacher complexity for $\beta$-mixing processes (e.g., Doukhan, 1994).
@@
The illustrative simulation studies in Section 5 show that the theoretical results derived in Sections 34 find empirical evidence in practice, including also in scenarios characterized by model misspecification and data contamination (theoretical and empirical results under non-i.i.d. data generating processes can be found in Appendix C of the Supplementary Material). These findings suggest that discrepancies with guarantees of uniformly-vanishing Rademacher complexity provide a robust and sensible choice when the assumed statistical model and/or the underlying data generating process do not necessarily meet specific regularity conditions, or it is not possible to verify these conditions. This is a common situation in applications, since the data generating process is unknown in practice.
@@
As discussed in Section 6, the unexplored bridge between discrepancy-based abc and the Rademacher complexity introduced in this article can also be leveraged to derive even more general theory by exploiting the active literature on the Rademacher complexity. For example, combining our perspective with the recent unified treatment of ips and $f$-divergences (Agrawal and Horel, 2021; Birrell et al., 2022) might set the premises to derive similarly-interpretable and general results for other discrepancies employed within abc, such as the Kullback–Leibler divergence (Jiang, Wu and Wong, 2018) and Hellinger distance (Frazier, 2020). More generally, our contribution can have implications even beyond abc, in particular on generalized Bayesian inference via pseudo-posteriors based on discrepancies (Bissiri, Holmes and Walker, 2016; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022; Frazier, Knoblauch and Drovandi, 2024). Proofs and additional results can be found in the Supplementary Material.
2 Integral probability semimetrics and Rademacher complexity
@@
Denote with $y_{1:n}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$ an i.i.d. sample from $\mu^{*}\in\mathcal{P}(\mathcal{Y})$, where $\mathcal{P}(\mathcal{Y})$ is the space of probability measures on $\mathcal{Y}$, and assume that $\mathcal{Y}$ is a metric space endowed with distance $\rho$; see Appendix C in the Supplementary Material for extensions to non-i.i.d. settings. Given a statistical model $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ in $\mathcal{P}(\mathcal{Y})$ and a prior distribution $\pi$ on $\theta$, rejection abc iteratively samples $\theta$ from $\pi$, draws a synthetic i.i.d. sample $z_{1:m}=(z_{1},\ldots,z_{m})$ from $\mu_{\theta}$, and retains $\theta$ as a sample from the abc posterior if the discrepancy $\Delta(z_{1:m},y_{1:n})$ between $z_{1:m}$ and $y_{1:n}$ is below a user-selected threshold $\varepsilon_{n}\geq 0$. Although $m$ may differ from $n$, here we follow common practice in abc theory (e.g., Bernton et al., 2019; Frazier, 2020) and set ${m=n}$. Rejection abc does not sample from the exact posterior $\pi_{n}(\theta)\propto\pi(\theta)\mu_{\theta}^{n}(y_{1:n}),$ but rather from the abc posterior defined as
@@
whose properties clearly depend on the chosen discrepancy $\Delta(\cdot,\cdot)$. Within summary-based abc, $\Delta(z_{1:n},y_{1:n})$ is a suitable distance, typically Euclidean, among summaries computed from the synthetic sample $z_{1:n}$ and the observed data $y_{1:n}$. However, recalling, e.g., Marin et al. (2014) and Frazier et al. (2018), the identification of summaries that do not lead to information loss is challenging for those complex models requiring abc implementations.
@@
where $\delta_{x}$ is the Dirac delta at a generic $x\in\mathcal{Y}$. Popular examples are abc versions based on mmd, kl divergence, Wasserstein distance, energy statistic, Hellinger and Cramer–von Mises distances, and $\gamma$-divergence, whose limiting properties have been studied in Park, Jitkrittum and Sejdinovic (2016), Jiang, Wu and Wong (2018), Bernton et al. (2019), Nguyen et al. (2020), Frazier (2020) and Fujisawa et al. (2021) under different asymptotic regimes and with a common reliance on existence assumptions to ease the proofs. As a first step toward our unified and constructive theoretical framework, we shall emphasize that, although most of the above contributions treat the discrepancies separately, some of these choices share, in fact, a common origin. For example, mmd, Wasserstein distance and the energy statistic all belong to the ips class (Müller, 1997) in Definition 2.1. Such a class also includes summary-based distances.
Definition 2.1 (Integral probability semimetric — ips).
Let $\mathfrak{F}$ be a class of measurable functions $f:\mathcal{Y}\to\mathds{R}$. Then, the integral probability semimetric $\mathcal{D}_{\mathfrak{F}}$ between $\mu_{1},\mu_{2}\in\mathcal{P}(\mathcal{Y})$ is defined as
@@
where $\mathfrak{F}$ is the pre-specified class of functions that characterizes $\mathcal{D}_{\mathfrak{F}}$.
Examples 2.22.4 show that routinely-employed discrepancies both in summary-free and summary-based abc (e.g., Park, Jitkrittum and Sejdinovic, 2016; Bernton et al., 2019; Nguyen et al., 2020; Drovandi and Frazier, 2022) are, in fact, ips with a known characterizing family $\mathfrak{F}$ that uniquely identifies each discrepancy.
@@
(Wasserstein-1 distance). Let us consider the Lipschitz seminorm defined as $||f||_{L}:=\sup\{|f(x)-f(x^{\prime})|/\rho(x,x^{\prime}):x\neq x^{\prime}\mbox{ in }\mathcal{Y}\}$. If $\mathfrak{F}=\{f:||f||_{L}\leq 1\}$, then $\mathcal{D}_{\mathfrak{F}}$ coincides with the Kantorovich metric. Recalling the Kantorovich–Rubinstein theorem (e.g., Dudley, 2018), when $\mathcal{Y}$ is separable this metric coincides with the dual representation of the Wasserstein-1 distance, which is therefore recovered as an example of ips. Recall also that such a distance is defined on the space of probability measures having finite first moment (e.g., Sriperumbudur et al., 2012; Bernton et al., 2019).
@@
(mmd and energy distance). Given a positive-definite kernel $k(\cdot,\cdot)$ on $\mathcal{Y}\times\mathcal{Y}$, let $\mathfrak{F}=\{f:||f||_{\mathcal{H}}\leq 1\}$, where $\mathcal{H}$ is the reproducing kernel Hilbert space corresponding to $k(\cdot,\cdot)$. Then $\mathcal{D}_{\mathfrak{F}}$ is the mmd (e.g., Muandet et al., 2017). When $k(\cdot,\cdot)$ is characteristic, then $\mathcal{D}_{\mathfrak{F}}$ is not only a semimetric, but also a metric — i.e., $\mathcal{D}_{\mathfrak{F}}(\mu_{1},\mu_{2})=0$ implies $\mu_{1}=\mu_{2}$. Relevant examples of routinely-implemented characteristic kernels in $\mathcal{Y}=\mathbb{R}^{d}$ are the Gaussian $\exp(-\|x-x^{\prime}\|^{2}/\sigma^{2})$ and Laplace $\exp(-\|x-x^{\prime}\|/\sigma)$ ones (Muandet et al., 2017). Note that mmd is also directly related to the energy distance, due to the correspondence among positive-definite kernels and negative-definite functions (Sejdinovic et al., 2013).